OpenStudy (anonymous):

Write the equation of a circle whose center is at the origin and that contains the point (-3, 0).

jimthompson5910 (jim_thompson5910):

Hint: The general equation of any circle is (x-h)^2+(y-k)^2=r^2

jimthompson5910 (jim_thompson5910):

where (h,k) is the center and r is the radius

jimthompson5910 (jim_thompson5910):

Keep in mind that the origin is the point (0,0). So to find the radius, you need to find the distance from (0,0) to (-3,0)

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so r = 3

jimthompson5910 (jim_thompson5910):

the center is at the origin, so (h,k) =(0,0) which means h = 0 and k = 0

OpenStudy (anonymous):

and x>?

jimthompson5910 (jim_thompson5910):

I'm not sure what you mean, but the equation (x-h)^2+(y-k)^2=r^2 becomes (x-0)^2+(y-0)^2=3^2 x^2 + y^2 = 9